Calculate the pH for each of the cases in the titration of 25.0 mL of 0.200 M pyridine, C5H5N(aq) with 0.200 M HBr(aq).

a)  before addition of any HBr

b) after addition of 12.5 mL of HBr

c)  after addition of 15.0 mL of HBr

d) after addition of 25.0 mL of HBr

e) after addition of 29.0 mL of HBr
Answer #1

Ans:

Pyridine is a weak base.

Kb of pyridine = 1.7 x 10-9(data collected)

pKb of pyridine = -log Kb = 8.77

Reaction: C5H5N + HBr --> (C5H5NH)+ Br- (pyridinium bromide)

Taken

Taken (converted to millimole)

Millimole = Molarity x volume (mL)

What happens?

Remark

a

C5H5N = 25.0 mL, 0.20 M

C5H5N = 5 mmol

Weak base

b

C5H5N = 25.0 mL, 0.20 M

HBr = 12.5 mL, 0.20 M

C5H5N = 5 mmol

HBr = 2.5 mmol

2.5 mmol of HBr reacts with 2.5 mmol of pyridine to give 2.5 mmol of pyridinium bromide.

So, after the reaction we will have

Pyridine = 2.5 mmol

Pyridinium bromide = 2.5 mmol

BASIC BUFFER

c

C5H5N = 25.0 mL, 0.20 M

HBr = 15.0 mL, 0.20 M

C5H5N = 5 mmol

HBr = 3 mmol

3.0 mmol of HBr reacts with 3.0 mmol of pyridine to give 3.0 mmol of pyridinium bromide.

So, after the reaction we will have

Pyridine = 2.0 mmol

Pyridinium bromide = 3.0 mmol

BASIC BUFFER

d

C5H5N = 25.0 mL, 0.20 M

HBr = 25.0 mL, 0.20 M

C5H5N = 5 mmol

HBr = 5 mmol

5.0 mmol of HBr reacts with 5.0 mmol of pyridine to give 5.0 mmol of pyridinium bromide.

So, after the reaction we will have

Pyridine = 0.0 mmol

Pyridinium bromide = 5.0 mmol

SALT

e

C5H5N = 25.0 mL, 0.20 M

HBr = 29.0 mL, 0.20 M

C5H5N = 5 mmol

HBr = 5.8 mmol

5.0 mmol of HBr reacts with 5.0 mmol of pyridine to give 5.0 mmol of pyridinium bromide.

So, after the reaction we will have

Pyridine = 0.0 mmol

Pyridinium bromide = 5.0 mmol

HBr = 0.8 mmol

SALT + ACID

After the reaction

Volume of the solution = Volume of pyridine + volume of HBr

New Molarity

= mmol/volume(mL)

Remark

a

Pyridine = 5.0 mmol

25.0 mL

0.2 M

Weak base

b

Pyridine = 2.5 mmol

Pyridinium bromide = 2.5 mmol

37.5 mL

Pyridine = 0.067 M

Pyridinium bromide = 0.067

BASIC BUFFER

c

Pyridine = 2.0 mmol

Pyridinium bromide = 3.0 mmol

40.0 mL

Pyridine = 0.05 M

Pyridinium bromide = 0.075 M

BASIC BUFFER

d

Pyridine = 0.0 mmol

Pyridinium bromide = 5.0 mmol

50.0 mL

Pyridinium bromide = 0.1 M

SALT

e

Pyridine = 0.0 mmol

Pyridinium bromide = 5.0 mmol

HBr = 0.8 mmol

54.0 mL

Pyridinium bromide = 0.1 M

HBr = 0.015 M

SALT + ACID

Kb of pyridine = 1.7 x 10-9 (data collected)

pKb of pyridine = -log Kb = 8.77

New Molarity

= mmol/volume(mL)

Remark

pH

a

Pyridine = 0.2 M (C)

Weak base

[OH-] = (Kb. C)½ = (1.7 x 10-9 x 0.2)½ = 1.844 x 10-5 M

pOH = -log[OH-] = 4.734

pH = 14-pOH = 9.266

b

Pyridine = 0.067 M

Pyridinium bromide = 0.067

BASIC BUFFER

Henderson’s equation

pH = 14-pOH = 5.23

c

Pyridine = 0.05 M

Pyridinium bromide = 0.075 M

BASIC BUFFER

Henderson’s equation

pH = 14-pOH = 5.054

d

Pyridinium bromide = 0.1 M (C)

SALT

It is a salt made up of weak base and strong acid.

e

Pyridinium bromide = 0.1 M

HBr = 0.015 M (C)

SALT + ACID

Though we have salt and acid, contribution of salt to the pH is insignificant in comparison to the acid contribution. So pH will be majorly decided by acid.

[H+] = C

pH = -log[H+] = -log(0.015)= 1.824